Following are two boxes containing colored and white balls. A ball is drawn at random from box 1. Then a ball is drawn at random from box 2, and the colors of balls from both boxes are recorded in order. Find each of the following:
a. The probability of two white balls
b. The probability of at least one colored ball
c. The probability of at most one colored ball
d. The probability of colored ball white ball or white ball colored ball
My answers are:
a. 1/5
b. 4/5
c. 7/10
d. 1/2
Could you please explain why my answers are incorrect really trying to get a better understanding.

Answer by stanbon(75887)

Following are two boxes containing colored and white balls. A ball is drawn at random from box 1. Then a ball is drawn at random from box 2, and the colors of balls from both boxes are recorded in order. Find each of the following:
--------
Box 1: 2 Black Balls 3White Balls.
Box 2: 4 Black Balls 2 White Balls.
----------
a. The probability of two white balls::(3/5)(2/6) = 1/5
---------------------------
b. The probability of at least one colored ball::: 1-[1/5] = 4/5
---------------------------
c. The probability of at most one colored ball::
P(0 or 1 colored ball) = 1-P(2 colored) = 1-(2/5)(4/6) = 1-(8/30) = 22/30
= 11/15
-----------------------------
d. The probability of colored ball white ball or white ball colored ball
(2/5)(2/6) + (3/5)(4/6) = 16/30 = 8/15
================
Cheers,
Stan H.
===========
My answers are:
a. 1/5
b. 4/5
c. 7/10
d. 1/2





You did not post the number of black or of white balls.
Cheers,
Stan H.